OpenStudy (anonymous):

Fool's problem of the day, If \(f(x,y+1) = f(x,y)+x+1\) and \(f(y,0) =y \forall x,y \in \mathbb{Z} \). Then can you find \(f(7,6)\)?

OpenStudy (anonymous):

PS: This is a very easy problem, I am uploading on request of @Callisto.

OpenStudy (anonymous):

Well, I gave you a headline for it ;)

OpenStudy (callisto):

Nah.. Thanks.. :P

OpenStudy (anonymous):

Haha, Thanks CS :D , now I wish to have posted something harder ;)

OpenStudy (experimentx):

f(7,0) = 7 f(7,1) = 7+8 = 16 f(7,2) = 16 + 8 .. ap??

OpenStudy (experimentx):

f(7,6) = 7+6*8 = 55??

OpenStudy (anonymous):

Yes, \[ f(7, 6) = f(7, 5) + 7 + 1 = f(7, 5) + 8 \] \[= f(7, 4) + 16 = f(7, 3) + 24\] \[ = f(7, 2) + 32 = f(7, 1) + 40 \] \[ = f(7, 0) + 48 = 7 + 48 = 55 \]

OpenStudy (anonymous):

f(x,y+1)=f(x,y)+x+1 f(7,6) = f(7,5) +7 +1 =f(7,4)+7+1+7+1 = f(7,4)+2(7)+2(1) = f(7,3)+7+1+7+1+7+1 = f(7,3)+3(7)+3(1) : : = f(7,0) +6(7)+6(1) = 7+42+6 =55

OpenStudy (anonymous):

I think I will post something harder now ;)

OpenStudy (anonymous):

Okay somewhat interesting : http://openstudy.com/study#/updates/4f746e7ae4b0b478589db2d5